OpenStudy (kl0723):

Integration of:

OpenStudy (kl0723):

\[\int\limits_{1}^{2} \frac{ e^{\frac{ 1 }{ x }} }{ x^{2} }dx\]

OpenStudy (loser66):

let u = 1/x and you can get the answer

OpenStudy (loser66):

don't forget changing the limits

OpenStudy (kl0723):

I'll give it a try, but what about the x on the bottom? and yes I'll change the limits once I'm on track

OpenStudy (loser66):

u =1/x --> du = -(1/2x^2) dx , that is \(-2du = \dfrac{1}{x^2} dx\) , right? no more denominator. :)

OpenStudy (kl0723):

thank you :)